OpenStudy (anonymous):

graph the linear equation Y-3=-2(x-2)

OpenStudy (anonymous):

y-3=-2(x-2) y-3=-2x+4 y=-2x+7 so -2 is the slope (=-2/1 so go down 2 and right 1) and the y-intercept is 7. now graph